OpenStudy (ell):

Solve the equation by using the quadratic formula: 4b^2 + 8b = 4

OpenStudy (anonymous):

first, get the equation into standard form: \(\ ax^2+bx+c=0 \), where a, b, c are real #s can you do that for me?

OpenStudy (ell):

4x^2 + 8x + 3 = 0...?

OpenStudy (anonymous):

where did that "+3" come from? check your algebra...

OpenStudy (ell):

haha(: sorry, i mixed 2 prob. up. My fault:P + 4 instead

OpenStudy (anonymous):

check it again... shouldn't it be " - 4" ??? so \(\ 4b^2 + 8b -4 = 0 \) ok???

OpenStudy (ell):

mmkay, b/c of the positive and negative... got it. Thanks!

OpenStudy (kenljw):

4b^2 + 8b + 4 = 0 b^2 + 2b + 1= 0 quadratic equation x^2 + bx +c = 0 x = (-b/2)(1 +- sqrt(1-c(2/b)^2) Therefore b=-1(1+-sqrt(0) = -1, -1 (b+1)^2 = 0

OpenStudy (ell):

yeah, ive got it now:) Thanks everyone for all your help@ Much appreciated! GBU!! <3 :)

OpenStudy (anonymous):

ok... so can you identify these for me? a = ??? b = ??? c = ???

OpenStudy (kenljw):

yes my form is used in engineering